OpenStudy (anonymous):

Describe the relationship between an equation in logarithmic form and an equivalent equation in exponential form.

OpenStudy (anonymous):

log function is the inverse function of exponential function

OpenStudy (anonymous):

The exponential function may be written as: y = bx. The exponential function is a one-to-one function, which means that for each x there is only one y and for each y there is only one x. Functions that are one-to-one have inverse functions. The relationship between a function and its inverse is that one function is the reflecion (about the line y = x) of the other. To write the equation of the inverse of a one-to-one function one must interchange the x and y variables and solve for y. For example, we can find the inverse function of y = 3x by first interchanging x and y (x = 3y) and then solving for y: y =1/3 x. Or we can consider the function y = x2 , x > 0, which has the inverse function x = y2 => y = . So y = bx has the inverse x = by

OpenStudy (anonymous):

no, exponential function is like this :3^x,4^x, y=bx is linear function

OpenStudy (anonymous):

ex log(base b)x=y raising both sides of powers b^[log(base b)x]=b^y x=b^y

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

lol..have fun guys